# Attitudes Towards and Knowledge of Non-Consensual Synthetic Intimate Imagery in 10 Countries

### 26 Jan 2024 v1

@@ Abstract.
Deepfake technology tools have become ubiquitous, ”democratizing” the ability to manipulate images and videos. One popular use of such technology is the creation of sexually explicit content, which can then be posted and shared widely on the internet. This article examines attitudes and behaviors related to non-consensual synthetic intimate imagery (NSII) across over 16,000 respondents in 10 countries. Despite nascent societal awareness of NSII, NSII behaviors were considered harmful. In regards to prevalence, 2.2% of all respondents indicated personal victimization, and 1.8% all of respondents indicated perpetration behaviors. Respondents from countries with relevant legislation also reported perpetration and victimization experiences, suggesting legislative action alone is not a sufficient solution to deter perpetration. Technical considerations to reduce harms may include suggestions for how individuals can better monitor their presence online, as well as enforced platform policies which ban, or allow for removal of, NSII content.
deepfake, pornography, image-based abuse, non-consensual explicit imagery, involuntary synthetic pornographic imagery
copyright: none
@@ 1. Introduction
Digitally altered content is a long-standing phenomenon, traditionally created through the use of photo-editing software to stitch together segments, add filters, or change the speed of videos. These older forms have now been dubbed ”cheapfakes” in contrast to ”deepfakes,” a portmanteau of ”deep learning” and ”fake” that describes AI-generated highly realistic content (Shahid et al., 2022; Westerlund, 2019). While deepfakes can be used in beneficial ways for accessibility and creativity (Farid, 2022; Cover, 2022), deepfakes can also be weaponized and used for malicious purposes, including financial fraud, disinformation dissemination, and sexual extortion (”sextortion”) (Farid, 2022; Aliman et al., 2021). Deepfake videos can be categorized into a variety of different types, including audio deepfakes, lip-syncing, puppet-master, face synthesis and attribute manipulation, and face-swapping (Wodajo and Atnafu, 2021). The focus of this paper is on face-swapping in the creation of deepfake pornography. This involves the face of the source person being replaced with the target person, so that the target person appears to engage in scenarios in which they never appeared (Masood et al., 2023). The terminology and language used has evolved in the image-based sexual abuse space more broadly (McGlynn et al., 2017) and accordingly, it is important to emphasize that deepfake pornography is, by its nature, involuntary and non-consensual. Therefore, throughout the paper, we use the term ”non-consensual synthetic intimate imagery” (NSII). One exception is when describing the language we used in the survey (where we used the more recognized ”deepfake pornography” term).
@@
The consumer creation of deepfakes started in late 2017 on Reddit, after a user named ”deepfakes” posted NSII depicting the faces of female celebrities superimposed onto pornographic videos on his page (Kugler and Pace, 2021; Westerlund, 2019). Continued consumer interest in NSII is reflected in the proliferation of dedicated NSII sites, often depicting celebrity targets. Abuse potential has increased in recent years as the technology has advanced in sophistication and availability (Cao et al., 2023; Masood et al., 2023; Gieseke, 2020; Wang, 2022). NSII can be considered a form of image-based sexual abuse because intimate images are created and/or shared without the consent of the person or persons depicted in the images. The harms of image-based sexual abuse have been well-documented, including negative impacts on victim-survivors’ mental health, career prospects, and willingness to engage with others both online and offline (Henry et al., 2020; Citron and Franks, 2014). The proliferation of NSII technologies means that anyone can now become a victim of image-based sexual abuse, and research suggests that women continue to bear the brunt of this online abuse (Kugler and Pace, 2021; Dunn, 2021). Investigating the prevalence of the misuse of deepfake technologies, specifically for NSII, can help Human Computer Interaction (HCI) research communities to better understand how best to mitigate the gendered societal harms associated with this phenomenon. Potential avenues for reduced harm may include comprehensive and considered legislation, NSII policies around the creation and/or distribution of content (e.g., Reddit’s continued de-platforming of subreddits dedicated to deepfake pornography or Google’s policy for removal of NSII), technical tools to help victim-survivors discover and automate takedown requests, and user experience treatments on deepfake creation tools designed to deter users from creating and distributing NSII.
@@
To begin mapping the prevalence of and general sentiment around NSII, we addressed the following research questions:
  • RQ1: What is the general public’s awareness of, and attitudes towards, NSII?
@@
  • RQ2: What is the prevalence of NSII behaviors (e.g., creating, viewing, and or/sharing)?
@@
In this study, we surveyed over 16,000 respondents in 10 different countries. Countries were selected based on a number of considerations, including representation of the diverse approach to legislation on the issue of image-based sexual abuse, replication/expansion of previous findings (e.g., in Australia and the United States) around the prevalence of image-based sexual abuse, and geographic diversity. Questions were designed to assess attitudes towards, and experiences of, image-based sexual abuse, with specific questions relating to NSII. This study was primarily quantitative, although optional open-ended questions allowed respondents to add additional context and thoughts. This research contributes to the existing literature by identifying trends in geographical awareness and behaviors; building on the existing image-based sexual abuse literature in terms of understanding who is most at-risk to this phenomenon; and probing the efficacy of current or proposed remedies.
@@ 2. Related Work
In this section, we provide a brief summary of the literature surrounding public perceptions of deepfakes more broadly and in relation to associated harms, deepfake detection, and legislative redress. The structure of this background review loosely follows that of our research questions; however, as prevalence data is a current gap in the literature, in our review below we also focus on whether existing or proposed laws sufficiently address the harms associated with this phenomenon.
@@ 2.1. Attitudes Towards Deepfakes
The implications of NSII both resemble, and are distinct to, traditional forms of image-based sexual abuse. In introducing the ”pervert’s dilemma,” Ohman (Öhman, 2020) notes that the mere creation of deepfake pornography in isolation is somewhat comparable to private sexual fantasies, which are not considered morally objectionable. However, when situated appropriately against the backdrop of gender inequality and gender-based violence, the moral unacceptability of NSII becomes clear. Moreover, deepfake technology provides a unique avenue for harassment and harm as it significantly reduces the barriers to the creation of non-consensual content (Akerley, 2020). As Harris (Harris, 2021) notes, while some have argued that concerns about the epistemic effects of deepfakes are overblown (e.g., because content is unlikely to be taken at face value and consumers will be able to critically assess veracity claims), artificially generated associations can have long-term harmful effects (see also (Rini and Cohen, 2022)).
@@
Researchers have begun to explore the different perspectives on deepfakes through attitudinal research. Although this body of work is still somewhat nascent, results suggest significant concern about the discovery and propagation of deepfake videos (Gottfried, 2019; Cochran and Napshin, 2021). Additionally, studies have focused on the harms associated with deepfakes, including a study published in 2021 in which respondents deemed creating and posting NSII, even when labeled as fake, as highly harmful and deserving of punishment (Kugler and Pace, 2021). One study (Fido et al., 2022) found that while female respondents indicated greater levels of victim harm, men were more likely to create NSII. The study also found that harms are greater when content is shared, as opposed to when it is used for personal sexual gratification. To the best of our knowledge, a study with Indian respondents is the sole general population survey outside of the United States. They found that awareness of deepfakes among Indian respondents was low, and even those who knew about deepfakes expected (incorrectly) to be able to easily discern fake videos (Shahid et al., 2022).
@@
A subset of research has also been conducted with respondents who are knowledgeable about deepfakes. These studies have focused primarily on ethical or privacy concerns. One study examined Reddit users actively discussing deepfakes, which found, via content analysis, that users were generally pro-deepfake technology, regardless of the consequences (Gamage et al., 2022). In contrast, qualitative research involving individuals using an open-source deepfake creation tool found that participants had reflexively engaged with the concepts of consent and privacy, and expressed significant concerns about the misuse of deepfakes (Widder et al., 2022). These themes were echoed and expanded upon in a mixed-methods study carried out in China (Li and Wan, 2023), where users familiar with deepfakes highlighted informed consent, privacy protection, and non-deception as relevant factors influencing the social acceptance of deepfakes. That is, if deepfakes are transparent and created or distributed with the knowledge and consent of those featured, they are more acceptable than deepfakes created non-consensually and/or for deceptive purposes. In addition to these factors, attitudes are also informed by the content or aim of the deepfakes (e.g., for creative as opposed to malicious purposes). For example, two separate studies found that entertainment (Li and Wan, 2023) or humorous/amusement value (Cochran and Napshin, 2021) mitigated ethical judgments on deepfakes. This may suggest that content such as NSII or parody videos could be considered more acceptable than videos intended to further political disinformation, for example.
@@
In sum, there is some quantitative, survey-based research suggesting that average respondents have concerns about the deleterious potential of deepfakes. Similarly, existing qualitative research with participants who have expertise on deepfakes shows that they share concerns about risks and harms. More research, however, is needed to investigate geographic diversity in general attitudinal quantitative research, as well as prevalence data, to broadly understand deepfake behavior. More work is also needed specifically on attitudes towards creating, sharing, or explicitly seeking out NSII.
@@ 2.2. Deepfake Creation and Detection
While researchers and journalists have been documenting the phenomenon of deepfakes for several years, interest has accelerated in recent years due to technological advancements that facilitate the creation of NSII in affordable and low-friction ways (Bond, 2023). For example, ”deepfake pornography,” as a keyword search within Google Scholar, returns 284 results from 2017-2019 as compared to 1,480 results from 2022-2023. Much of the research in this space (Gamage et al., 2021) focuses on documenting how this technology works or highlighting the performance issues for various detection tools or methods (Pu et al., 2021; Dolhansky et al., 2020; Masood et al., 2023; Stroebel et al., 2023; Rana and Sung, 2023). Like other general artificial intelligence research focusing on the issues associated with biased training data sets, deepfake detection research has identified diversity and bias problems within relevant datasets (Xu et al., 2022; Trinh and Liu, 2021). For example, detection techniques have been shown to be sensitive to gender, performing worse on female-based deepfakes (Nadimpalli and Rattani, 2022) as compared to male counterparts. A recent systematic review of deepfake detection papers succinctly concludes that current models are not robust to either novel or challenging conditions (Stroebel et al., 2023). This is particularly relevant to NSII, as non-digitally altered image-based sexual abuse material may be captured surreptitiously, or made with poor cameras and/or bad lighting conditions. NSII is thus likely to demonstrate the exact risk factors for poor detection rates.
In addition to technical tool failures, studies have documented the fallibility of humans in correctly identifying deepfakes as such (Köbis et al., 2021; Korshunov and Marcel, 2020; Thaw et al., 2020). This is seen across a variety of types of deepfakes, from deepfake audio (Mai et al., 2023) to images (Bray et al., 2023; Nightingale and Farid, 2022) and videos (Groh et al., 2022). It is possible that as deepfakes become more ubiquitous, humans will make improvements in identifying markers, however from a technological detection perspective, it seems likely that an arms race of sorts will develop, such that the increasing sophistication of deepfakes will require detection methods to make significant improvements to catch up and keep pace.
@@ 2.3. Deepfakes and the Law
The societal implications of deepfakes technology are far-reaching and, while many have been recognized or anticipated, current legislative remedies for misuse are typically deemed unlikely to be effective by experts (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). Although many scholars call for legislation, there are others who are skeptical of criminalizing NSII or deepfakes more broadly (Kirchengast, 2020; Feeney, 2021; Widder et al., 2022). Insufficient NSII laws should be considered the norm, rather than the exception, when considering the history of failed legislative proposals for image-based sexual abuse more broadly (Carter, 2022; Walker and Sleath, 2017; Akerley, 2020). In the legal sphere, the first American federal law regarding deepfakes was signed in December, 2019. This law focused on the potential of deepfakes to be weaponized by foreign powers to influence elections, and it also established a competition to encourage research into detection technologies (Ferraro et al., 2020). Four states (California, Virginia, New York, and Georgia) have also passed laws banning NSII specifically (Mania, 2020; Ryan-White et al., 2022; Kugler and Pace, 2021). At the end of 2022, a federal bill was introduced in the U.S. House of Representatives specifically prohibiting NSII (Morelle, 2023). Although legislative remedies have approached political deepfakes and NSII separately, one critique from scholars is that both operate in similar ways to silence critical speech (Maddocks, 2020).
Outside of the United States, several countries have either explicitly included NSII in existing laws or are considering making legislative amendments. For example, in Australia, some states and territories make it clear that the non-consensual distribution of intimate images includes digitally altered material as a criminal offense. The Australian Online Safety Act 2021 (of the eSafety Commissioner Australia, 2023) also provides a civil remedy scheme for a range of online harms, including the nonconsensual sharing of digitally altered images. South Korea specifically prohibits NSII, alongside all other forms of pornography (Lee, 2021) and in the England and Wales, the UK Online Safety Act 2023 criminalizes threatening or sharing intimate images without consent, including images that have been digitally altered (Lawson, 2023; eSafety Commissioner, nd). In most countries, however, and including most of those surveyed in our study (Belgium, Denmark, France, Mexico, Netherlands, Poland, Spain), there are no protections at all.

### 13 Feb 2024 v2

# Non-Consensual Synthetic Intimate Imagery: Prevalence, Attitudes, and Knowledge in 10 Countries

@@ Abstract.
Deepfake technologies have become ubiquitous, ”democratizing” the ability to manipulate photos and videos. One popular use of deepfake technology is the creation of sexually explicit content, which can then be posted and shared widely on the internet. Drawing on a survey of over 16,000 respondents in 10 different countries, this article examines attitudes and behaviors related to ”deepfake pornography” as a specific form of non-consensual synthetic intimate imagery (NSII). Our study found that deepfake pornography behaviors were considered harmful by respondents, despite nascent societal awareness. Regarding the prevalence of deepfake porn victimization and perpetration, 2.2% of all respondents indicated personal victimization, and 1.8% all of respondents indicated perpetration behaviors. Respondents from countries with specific legislation still reported perpetration and victimization experiences, suggesting NSII laws are inadequate to deter perpetration. Approaches to prevent and reduce harms may include digital literacy education, as well as enforced platform policies, practices, and tools which better detect, prevent, and respond to NSII content.
deepfake pornography, deepfakes, AI-generated images, image-based sexual abuse, non-consensual explicit imagery, involuntary synthetic pornographic imagery
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642382isbn: 979-8-4007-0330-0/24/05ccs: Social and professional topics Governmental regulationsccs: Applied computing Lawccs: Applied computing Psychologyccs: Applied computing Sociologyccs: Human-centered computing Human computer interaction (HCI)ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Non-consensual synthetic intimate imagery (NSII) refers to digitally altered content that is fake but depicts the faces, bodies, and/or voices of real people. NSII can be created through more traditional means using photo-editing software to stitch together segments, add filters, or change the speed of videos — often referred to as “shallowfakes” or “cheapfakes.” Increasingly, however, NSII is being created through the use of artificial intelligence (AI), involving different methods, such as speech-to-speech voice conversion, lip-syncing, puppet-master, face synthesis, attribute manipulation, and face-swapping (Wodajo and Atnafu, 2021). AI-generated NSII is more colloquially known as “deepfake pornography.” The consumer creation of deepfakes (a portmanteau of “deep learning” and “fake” (Shahid et al., 2022; Westerlund, 2019)) started in late 2017 on Reddit, after a user named “deepfakes” posted NSII depicting the faces of female celebrities “stitched” onto pornographic videos (Kugler and Pace, 2021; Westerlund, 2019). Continued consumer interest in deepfakes is reflected in the proliferation of dedicated deepfake sites and forums, often depicting celebrity targets. While deepfakes can be used in beneficial ways for accessibility and creativity (Farid, 2022; Cover, 2022), abuse potential has increased in recent years as the technology has advanced in sophistication and availability (Cao et al., 2023; Masood et al., 2023; Gieseke, 2020; Wang, 2022). Deepfakes can be weaponized and used for malicious purposes, including financial fraud, disinformation dissemination, cyberbullying, and sexual extortion (“sextortion”) (Farid, 2022; Aliman et al., 2021).
@@
Non-consensual deepfake pornography can be considered a form of image-based sexual abuse because intimate images are created and/or shared without the consent of the person or persons depicted in the images. The harms of image-based sexual abuse have been well-documented, including negative impacts on victim-survivors’ mental health, career prospects, and willingness to engage with others both online and offline (Henry et al., 2020; Citron and Franks, 2014). The proliferation of NSII technologies means that anyone can now become a victim of image-based sexual abuse, and research suggests that women continue to bear the brunt of this online abuse (Kugler and Pace, 2021; Dunn, 2021). Investigating the prevalence of the misuse of deepfake technologies, specifically for NSII, can help Human Computer Interaction (HCI) research communities to better understand how best to mitigate the gendered societal harms associated with this phenomenon. Potential avenues for reduced harm may include comprehensive and considered legislation, NSII policies around the creation and/or distribution of content (e.g., Reddit’s continued de-platforming of subreddits dedicated to deepfake pornography or Google’s policy for removal of NSII), technical tools to help victim-survivors discover and automate takedown requests, and user experience treatments on deepfake creation tools designed to deter users from creating and distributing NSII.
Over the past decade, the terminology and language has evolved in the image-based sexual abuse space to move beyond problematic, perpetrator terms likes “revenge pornography” to better encapsulate the non-consensual nature of such acts (McGlynn et al., 2017). Accordingly, it is important to emphasize that deepfake pornography is, by its nature, both involuntary and non-consensual. Therefore, throughout the article, we use the broader term “non-consensual synthetic intimate imagery” (NSII) to refer to fake, digitally altered images created using AI and non-AI tools, as well as the more specific term “AI-generated image-based sexual abuse” (AI-IBSA) to refer to fake, digitally altered images created using AI. However, when reporting on our survey questions and findings, we use the more recognized “deepfake pornography” term (as well as “deepfakes”) to avoid confusion for the reader. We focus on face-swapping which involves replacing the face of the source person with the target person, so that the target person appears to engage in scenarios in which they never appeared (Masood et al., 2023).
@@
To begin mapping the prevalence of and general sentiment around AI-IBSA, we addressed the following research questions:
  • RQ1: What is the general public’s awareness of, and attitudes towards, AI-IBSA?
@@
  • RQ2: What is the prevalence of AI-IBSA behaviors (e.g., creating, viewing, and or/sharing images)?
@@
In this study, we surveyed over 16,000 respondents in 10 different countries. Countries were selected based on a number of considerations, including representation of the diverse approach to legislation on the issue of image-based sexual abuse, replication/expansion of previous findings (e.g., in Australia and the United States) around the prevalence of image-based sexual abuse, and geographic diversity. Questions were designed to assess attitudes towards, and experiences of, image-based sexual abuse, with specific questions relating to NSII and more specifically AI-IBSA. The study was primarily quantitative, although optional open-ended questions allowed respondents to add additional context and thoughts. The research contributes to the existing literature by: identifying trends in geographical awareness and behaviors; building on the existing image-based sexual abuse literature in terms of understanding who is most at-risk to this phenomenon; and, probing the efficacy of current or proposed remedies.
@@ 2. Related Work
In this section, we provide a brief summary of the literature surrounding public perceptions of deepfakes more broadly and in relation to associated harms, deepfake detection, and legislative redress. The structure of this background review loosely follows that of our research questions; however, as prevalence data is a current gap in the literature, in our review below we also focus on whether existing or proposed laws sufficiently address the harms associated with AI-IBSA.
@@ 2.1. Attitudes Towards Deepfakes
AI-IBSA is both similar to and distinct from traditional forms of image-based sexual abuse. In introducing the “pervert’s dilemma,” Ohman (Öhman, 2020) notes that the mere creation of deepfake pornography in isolation is somewhat comparable to private sexual fantasies, which are not considered morally objectionable. However, when situated appropriately against the backdrop of gender inequality and gender-based violence, the moral unacceptability of AI-IBSA becomes clear. Moreover, deepfake technology provides a unique avenue for harassment and harm as it significantly reduces the barriers to the creation of non-consensual content (Akerley, 2020). As Harris (Harris, 2021) notes, while some have argued that concerns about the epistemic effects of deepfakes are overblown (e.g., because content is unlikely to be taken at face value and consumers will be able to critically assess veracity claims), artificially generated associations can have long-term harmful effects (see also (Rini and Cohen, 2022)).
@@
Researchers have begun to explore the different perspectives on deepfakes through attitudinal research. Although this body of work is still somewhat nascent, results suggest significant concern about the discovery and propagation of deepfake videos (Gottfried, 2019; Cochran and Napshin, 2021). Additionally, studies have focused on the harms associated with deepfakes, including a study published in 2021 in which respondents deemed creating and posting deepfake pornography, even when labeled as fake, as highly harmful and deserving of punishment (Kugler and Pace, 2021). One study (Fido et al., 2022) found that while female respondents indicated greater levels of victim harm, men were more likely to create deepfake pornography. The study also found that harms are greater when content is shared, as opposed to when it is used for personal sexual gratification. To the best of our knowledge, a study with Indian respondents is the sole general population survey outside of the United States. They found that awareness of deepfakes among Indian respondents was low, and even those who knew about deepfakes expected (incorrectly) to be able to easily discern fake videos (Shahid et al., 2022).
@@
A subset of research has also been conducted with respondents who are knowledgeable about deepfakes. These studies have focused primarily on ethical or privacy concerns. One study examined Reddit users actively discussing deepfakes, which found, via content analysis, that users were generally pro-deepfake technology, regardless of the consequences (Gamage et al., 2022). In contrast, qualitative research involving individuals using an open-source deepfake creation tool found that participants had reflexively engaged with the concepts of consent and privacy, and expressed significant concerns about the misuse of deepfakes (Widder et al., 2022). These themes were echoed and expanded upon in a mixed-methods study carried out in China (Li and Wan, 2023), where users familiar with deepfakes highlighted informed consent, privacy protection, and non-deception as relevant factors influencing the social acceptance of deepfakes. That is, if deepfakes are transparent and created or distributed with the knowledge and consent of those featured, they are more acceptable than deepfakes created non-consensually and/or for deceptive purposes. In addition to these factors, attitudes are also informed by the content or aim of the deepfakes (e.g., for creative as opposed to malicious purposes). For example, two separate studies found that entertainment (Li and Wan, 2023) or humorous/amusement value (Cochran and Napshin, 2021) mitigated ethical judgments on deepfakes. This may suggest that content such as sexual or parody videos could be considered more acceptable than videos intended to further political disinformation, for example.
@@
In sum, there is some quantitative, survey-based research suggesting that average respondents have concerns about the deleterious potential of deepfakes. Similarly, existing qualitative research with participants who have expertise on deepfakes shows that they share concerns about risks and harms. More research, however, is needed to investigate geographic diversity in general attitudinal quantitative research, as well as prevalence data, to broadly understand deepfake behavior and its impacts. More work is also needed specifically on attitudes towards creating, sharing, or explicitly seeking out such content.
@@ 2.2. Deepfake Creation and Detection
While researchers and journalists have been documenting the phenomenon of deepfakes for several years, interest has accelerated in recent years due to technological advancements that facilitate the creation of deepfakes in affordable and low-friction ways (Bond, 2023). For example, “deepfake pornography,” as a keyword search within Google Scholar, returns 284 results from 2017-2019 as compared to 1,480 results from 2022-2023. Much of the research in this space (Gamage et al., 2021) focuses on documenting how this technology works or highlighting the performance issues for various detection tools or methods (Pu et al., 2021; Dolhansky et al., 2020; Masood et al., 2023; Stroebel et al., 2023; Rana and Sung, 2023). Like other broader AI research focusing on the issues associated with biased training data sets, deepfake detection research has identified diversity and bias problems within relevant datasets (Xu et al., 2022; Trinh and Liu, 2021). For example, detection techniques have been shown to be sensitive to gender, performing worse on female-based deepfakes (Nadimpalli and Rattani, 2022) as compared to images of males. A recent systematic review of deepfake detection papers succinctly concludes that current models are sensitive to both novel or challenging conditions (Stroebel et al., 2023). This is particularly relevant, as non-digitally altered image-based sexual abuse material may be captured surreptitiously, or made with poor cameras and/or bad lighting conditions. AI-IBSA is thus likely to demonstrate the exact risk factors for poor detection rates.
In addition to technical tool failures, studies have documented the fallibility of humans in correctly identifying deepfakes as fake (Köbis et al., 2021; Korshunov and Marcel, 2020; Thaw et al., 2020). This is seen across a variety deepfake types, from deepfake audio (Mai et al., 2023) to photos (Bray et al., 2023; Nightingale and Farid, 2022) and videos (Groh et al., 2022). It is possible that as deepfakes become more ubiquitous, humans will improve in identifying markers or other signals suggesting the content is not “real.” However from a technological detection perspective, it seems likely that an arms race of sorts will develop, such that increasingly sophisticated deepfakes will require significant improvements in detection methods to catch up and keep pace.
@@ 2.3. Deepfakes and the Law
The societal implications of deepfakes technology are far-reaching and, while many have been recognized or anticipated, experts generally agree current legislative remedies for misuse are unlikely to be effective (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). Although many scholars call for legislation, there are others who are skeptical of criminalizing NSII, AI-IBSA, or deepfakes more broadly (Kirchengast, 2020; Feeney, 2021; Widder et al., 2022). Insufficient NSII laws should be considered the norm, rather than the exception, when considering the history of failed legislative proposals for image-based sexual abuse in general (Carter, 2022; Walker and Sleath, 2017; Akerley, 2020). In the legal sphere, the first American federal law regarding deepfakes was signed in 2019. This law focuses on the potential of deepfakes to be weaponized by foreign powers to influence elections, and it also establishes a competition to encourage research into detection technologies (Ferraro et al., 2020). Four states (California, Virginia, New York, and Georgia) have also passed laws criminalizing “deepfake pornography” specifically (Mania, 2020; Ryan-White, 2022; Kugler and Pace, 2021). At the end of 2022, a federal bill was introduced in the U.S. House of Representatives specifically prohibiting deepfake pornography (Morelle, 2023). Although legislative remedies have approached political deepfakes and AI-IBSA separately, one critique from scholars is that both operate in similar ways to chill critical speech (Maddocks, 2020).
Outside of the United States, several countries have either explicitly included NSII in existing laws or are considering making legislative amendments. For example, in Australia, some states and territories make it clear that the non-consensual distribution of intimate images includes digitally altered material as a criminal offense. The Australian Online Safety Act 2021 (of the eSafety Commissioner Australia, 2023) also provides a civil remedy scheme for a range of online harms, including the non-consensual sharing of digitally altered images. South Korea specifically prohibits NSII, alongside all other forms of pornography (Lee, 2021). And in England and Wales, the UK Online Safety Act 2023 criminalizes threatening or sharing intimate images without consent, including images that have been digitally altered (Lawson, 2023; eSafety Commissioner, 2023). In most countries, however, and including most of those surveyed in our study (Belgium, Denmark, France, Mexico, Netherlands, Poland, Spain), there are no protections at all.